OpenStudy (anonymous):

Am I doing this right? What is the sum of the geometric sequence -1, 6, -36, if there are 7 terms? First term: -1 Common ratio: -30 N: How do I find "n" again???

hartnn (hartnn):

n=number of terms =7

OpenStudy (anonymous):

oh really? okay perfect thanks!!!!

hartnn (hartnn):

common ratio =-30 ???? :O thats incorrect -_-

OpenStudy (anonymous):

Common ratio..;;what hartnn said!

OpenStudy (anonymous):

the answer i got was wrong....

hartnn (hartnn):

common ratio = 6/(-1) = -36/6=.... ?

OpenStudy (anonymous):

omg i used the steps for the arthmetic sequence!!!!!!!!!!!

OpenStudy (anonymous):

= -6

hartnn (hartnn):

know the formula for the sum of the geometric sequence or should i mention it ?

OpenStudy (anonymous):

No but I'll write it down

hartnn (hartnn):

r=-6 , a1 =-1, n=7 are correct.

OpenStudy (anonymous):

ok:)

OpenStudy (anonymous):

wait how did you get a1 = -1?

hartnn (hartnn):

a1= 1st term an =n'th term

OpenStudy (anonymous):

wait where does it state the n'th term??

OpenStudy (anonymous):

sorry for my foolish questions

hartnn (hartnn):

it doesn't , and you don't need it here. just an extra info. here.

OpenStudy (anonymous):

i got my answer!!!!!! thank you soooooooooo much

hartnn (hartnn):

welcome ^_^